OpenStudy (anonymous):

f(x)= 7x+5; g(x)=x^2. Perform function operation of (f+g)(x) and find the domain.

OpenStudy (shadowfiend):

First off, remember that \((f+g)(x) = f(x) + g(x)\). So what is \(f(x) + g(x)\) above?

OpenStudy (anonymous):

isn't it \[x^{2} +7x+5?\]

OpenStudy (shadowfiend):

Right-o.

OpenStudy (shadowfiend):

Now, what is the `domain' of a function?

OpenStudy (anonymous):

i read its the x-value.

OpenStudy (shadowfiend):

Right. So is there any value of x for which the equation above doesn't give an answer that makes sense?

OpenStudy (anonymous):

hmm?

OpenStudy (shadowfiend):

The range is all real numbers unless there is a value of x for which: \[y = x^2 + 7x + 5\] Gives an undefined y.

OpenStudy (shadowfiend):

(Hint: for this equation, there is no such value -- all numbers you plug in for x will give you a y value. So, the domain is all real numbers, or \((-\infty, \infty)\).

OpenStudy (anonymous):

so its all real numbers? so how do you know when an equation's domain is limited?

OpenStudy (shadowfiend):

Well, take the equation \(y = \sqrt{x}\).

OpenStudy (shadowfiend):

If \(x \lt 0\), then the square root is undefined -- it produces an imaginary number. So, the domain of the function is any value >= 0.

OpenStudy (anonymous):

i get it you can't have negatives in a square root. so its either a whole number or zero right?

OpenStudy (shadowfiend):

It's either a *positive* whole number or zero. Negatives are also whole numbers.

OpenStudy (anonymous):

okay i get it finally thanks a bunch!! you've been very helpful. :0]

OpenStudy (shadowfiend):

No problem, glad to help!